a block of mass of 8kg resting on a horizontal rough horizontal table is connected by a light inextensible string which passes over a light frictionless pulley to another block of mass 5kg


1
Expert's answer
2021-08-05T11:35:29-0400

Given:

"m_1=8\:\rm kg"

"m_2=5\:\rm kg"

The Newton's second law says

"(m_1+m_2)a=m_2g"

Hence, the acceleration of the system

"a=\frac{m_2g}{m_1+m_2}=\frac{5\:\rm kg*9.8\: m/s^2}{(8+5)\:\rm kg}=3.8\:\rm m/s^2"